OpenStudy (anonymous):

log (5 + x) - log (x - 2) = log 2

OpenStudy (anonymous):

I have no idea...

OpenStudy (anonymous):

i love that name...:)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

start with \[\log(\frac{x+5}{x-2})=\log(2)\] and go from there

OpenStudy (anonymous):

 can you take it from that last step?